# Adaptive Width Neural Networks

### v2

@@ 3.1 Imposing a Soft Ordering on Neurons’ Importance
Now that the learning objective has been formalized, the missing ingredient is the definition of the neural network $p(y_{i}|\boldsymbol{\lambda}\text{=}\boldsymbol{\nu},\boldsymbol{\theta}\text{=}\boldsymbol{\rho},x_{i})$. Compared to a standard MLP, we need to make use of the variational parameters $\boldsymbol{\nu}$ that affect the truncation width at each hidden layer, whereas $\boldsymbol{\rho}$ corresponds to the the usual MLP weights. We choose a monotonically decreasing function $f_{\ell}$ for the simple reason that when a new neuron is added, its relative importance is low and will not drastically impact the network. As a result, we are imposing a soft ordering of importance among neurons.
We modify the classical activation $h^{\ell}_{j}$ of a hidden neuron $j$ at layer $\ell$ as
@@
where $D_{\ell-1}$ is the truncated width of the previous layer, $\sigma$ is a non-linear activation function and $w^{\ell}_{jk}\in\rho_{\ell j}$. That is, we rescale the activation of each neuron $k$ by its “importance” $f_{\ell}(j;\nu_{\ell})$. Note that the bias parameter is taken into account by concatenating a dummy value 1 to $h^{\ell-1}_{k}$.
It is easy to see that, in theory, the optimization algorithm could rescale the weights of the next layer by a factor $1/f_{\ell}(j;\nu_{\ell})$ to compensate for the term $f_{\ell}(j;\nu_{\ell})$. This may lead to a degenerate situation, which is actually quite common in classical MLPs, where the activations of the first neurons are small relative to the others, thus breaking the soft-ordering and wasting neurons. There are two strategies to address this undesirable effect. The first is to regularize the magnitude of the weights thanks to the prior $p(\theta_{\ell+1n})$, so that it may be difficult to compensate for the least important neurons that have a high $1/f_{\ell}(j)$. The second and less obvious strategy is to prevent the units’ activations of the current layer to compensate for high values by bounding their range. We apply both strategies to our experiments.
@@ 3.2 Rescaled Weight Initialization for Deep AWNN
Rescaling the activations of hidden units using Equation 14 causes activations of deeper layers to quickly decay to zero, as shown in Figure 2 for a 5-layer AWNN MLP with ReLU nonlinearity initialized using the well known kaiming scheme (He et al., 2015). This affects convergence since gradients get close to zero and it becomes impractical to train deep AWNN MLPs. We therefore derive a rescaled kaiming weight that guarantees that the variance of activation across layers is constant at initialization.
@@
Let us consider an MLP with activations as in Equation 14 and ReLU nonlinearity. At initialization, given $\alpha_{j}^{\ell}=\sigma\left(\sum_{k=1}^{D_{\ell-1}}w^{\ell}_{jk}h^{\ell-1}_{k}\right)$
$\mathrm{Var}[w_{j*}^{\ell}]=\frac{2}{\sum_{j=1}^{D_{\ell-1}}f^{2}_{\ell}(j)}\Rightarrow\mathrm{Var}[\alpha_{j}^{\ell}]\approx\mathrm{Var}[\alpha_{j}^{\ell-1}]$ (15)
@@
See Appendix A for an extended proof. ∎
The extended proof shows there is a connection between the variance of activations and the variance of gradients. In particular, the sufficient conditions over $\mathrm{Var}[w_{j*}^{\ell}]$ are identical if one initializes all $\nu_{\ell}$ in the same way for all layers. Therefore, if we initialize weights from a Gaussian distribution with standard deviation $\frac{\sqrt{2}}{\sqrt{\sum_{j=1}^{D_{\ell-1}}f^{2}_{\ell}(j)}}$, we guarantee that at initialization the variance of the deep network’s gradients will be constant. The effect of the new initialization (dubbed “kaiming+”) can be seen in Figure 2, where the distribution of activation values at initialization does not collapse in subsequent layers. This change drastically impacts overall convergence on the SpiralHard synthetic dataset (described in Section 4), where it appears it would be otherwise hard to converge using a standard kaiming initialization (please refer to Appendix A for a visual example).
Figure 2: Effect of the rescaled initialization scheme (“kaiming+”) on a ReLU-based AWNN MLP, where neurons’ activations are computed using Equation 14. Compared to the standard initialization, the variance of activations agrees with the theoretical result.
3.3 Future Directions and Limitations
MLPs are ubiquitous in modern deep architectures. They are used at the very end of CNNs, in each encoder and decoder layer of Transformer, and they process messages coming from neighbors in DGNs. Our experiments focus on MLPs to showcase AWNN’s broad applicability, but there are many other scenarios where one can apply AWNN’s principles. For instance, one could impose a soft ordering of importance on CNNs’ filters at each layer, therefore learning the number of filters during training. Alternatively, one could investigate the application of AWNN to timeseries, since recurrent networks (Hochreiter & Schmidhuber, 1997) can be seen as a special case of an MLP with weight sharing.
From a more theoretical perspective, we believe one could draw connections between our technique and the Information Bottleneck principle (Tishby et al., 2000), which seeks maximally representative (i.e., performance) and compact representations (e.g., width). Finally, the analysis of different functions $f$ to soft-order neurons, for instance a power-law distribution or a sigmoid-like function, may have a different impact on convergence and performance and could be empirically investigated.
The current limitation of AWNN is the relative overhead to change the network at training time. A minimal and naïve implementation has up to 10x cost compared to a fixed network (this relative overhead may reduce for more compute-bound tasks). Besides optimizing our code, other optimizations are conceivable: i) by updating the width every $M$ minibatches rather than at every step; ii) by allocating a larger network and dynamically grow it only if needed; iii) by creating ad-hoc support of AWNN on deep learning libraries like Pytorch. In any case, this overhead is relatively small compared to the exponential size of the hyper-parameter space when cross-validating the number of neurons of a multi-layered architecture.
@@ 3.3 Future Directions and Limitations
13:       $D_{\ell}\leftarrow\text{quantile function of }f_{\ell}(\cdot;\nu_{\ell})\text{\ evaluated at }k$
14:       Use $D_{\ell}$ to update $\boldsymbol{\rho}_{\ell},\boldsymbol{\rho}_{\ell+1}$ // add/remove neurons
15:    end for
16:  end function
@@ 4 Experiments and Setup
The purpose of the empirical analysis is not to claim AWNN is generally better than the fixed-width baseline. Rather, we demonstrate how AWNN overcomes the problem of fixing the number of neurons by learning it, thus reducing the amount of configurations to test. As such, due to the nature of this work and similarly to Mitchell et al. (2023), we use the remaining space to thoroughly study the behavior of AWNN, so that it becomes clear how to use it in practice. We first quantitatively verify that AWNN does not harm the performance compared to baseline models and compare the chosen width by means of grid-search model selection with the learned width of AWNN. Secondly, we check that AWNN chooses a higher width for harder tasks, which can be seen as increasing the hypotheses space until the neural network finds a good path to convergence. Third, we verify that convergence speed is not significantly altered by AWNN, so that the main limitation lies in the extra overhead for adapting the network at each training step. As a sanity check, we study conditions under which AWNN’s learned width does not seem to depend on starting hyper-parameters, so that their choice does not matter much. Finally, we analyze other practical advantages of training a neural network under the AWNN framework: the ability to compress information during training or post training, and the resulting trade-offs. Further analyses are in the Appendix.
@@
We now provide the details to reproduce our experiments on 9 datasets. We compare a baseline that undergoes proper hyper-parameter tuning against its AWNN version, where we replace any fixed MLP with an adaptive one. First, we train an MLP on 3 synthetic tabular tasks of increasing binary classification difficulty, namely a double moon, a spiral, and a double spiral that we call SpiralHard. A stratified hold-out split of 70% training/10% validation/20% test for risk assessment is chosen at random for these datasets. Similarly, we consider a ResNet-20 (He et al., 2016) trained on 3 image classification tasks, namely MNIST (LeCun, 1998), CIFAR10, and CIFAR100 (Krizhevsky, 2009), where data splits and preprocessing are taken from the original paper and AWNN is applied to the downstream classifier. In the graph domain, we train a Graph Isomorphism Network (Xu et al., 2019) on the NCI1 and REDDIT-B classification tasks using the same split and evaluation setup of Errica et al. (2020). These tasks are chosen as the structure was found to have a proper influence on the final performance. Here, the first 1 hidden layer MLP as well as the one used in each graph convolutional layer are replaced by adaptive AWNN versions. On all these tasks, the metric of interest is the accuracy. Finally, for the textual domain we train a Transformer architecture (Vaswani et al., 2017) on the Multi30k English-German translation task (Elliott et al., 2016), using a pretrained GPT-2 Tokenizer, and we evaluate the cross-entropy loss over the translated words. On tabular, image, and text-based tasks, an internal validation set (10%) for model selection is extracted from the union of outer training and validation sets, and the best configuration chosen according to the internal validation set is retrained 10 times on the outer train/validation/test splits, averaging test performances. Due to space reasons, we report datasets statistics and the hyper-parameter tried for the fixed and AWNN versions in Appendix B and C, respectively.
@@ 5 Results
We begin by discussing the quantitative results of our experiments: Table 1 reports means and standard deviations across the 10 final training runs. In terms of performance, we observe that AWNN is more stable or accurate than a fixed MLP on DoubleMoon, Spiral, and SpiralHard; all other things being equal, it seems that using more neurons and their soft ordering are the main contributing factors to these improvements. On the image datasets, performances of AWNN are comparable to those of the fixed baseline but for CIFAR100, due to an unlucky run that did not converge. In this case, AWNN learns a smaller total width compared to grid search.
Table 1: We compare performances and total width of MLP layers for the fixed and AWNN versions of the various models used. The exact width chosen by model selection on the graph datasets is unknown since we report published results. “Linear” means that the chosen downstream classifier is a linear model.
@@
Results on graph datasets are interesting in two respects: First, the performance on REDDIT-B is significantly improved by AWNN both in terms of average performance and stability of results; second, the total learned width is significantly higher than those tried in Xu et al. (2019); Errica et al. (2020), meaning that a biased choice of a good range of width has had a profound influence on the estimation of the risk for a specific family of DGN models (i.e., GIN). This result makes it evident that it is important to let the network decide how many neurons are necessary to solve the task. To check that the learned width is indeed conducive to good performances, Appendix D shows what happens when we retrain some fixed baselines using the total width as the width of each layer.
Finally, the results on the Multi30k show that the AWNN Transformer learns to use 200x parameters less than the fixed Transformer for the feed-forward networks, achieving a statistically comparable test loss. This result is appealing when read through the lenses of modern deep learning, as the power required by some neural networks such as Large Language Models (Brown et al., 2020) is so high that cannot be afforded by most institutions, and it demands future investigations.
5.1 Adaptation to Task Difficulty and Convergence
Intuitively, one would expect that AWNN learned larger widths for more difficult tasks. This is indeed what happens on the tabular datasets (and image datasets, see Appendix E) where some tasks are clearly harder than others. Figure 3 (left) shows that, given the same starting width per layer, the learned number of neurons grows according to the task’s difficulty. It is also interesting to observe that the total width for a multi-layer MLP on SpiralHard is significantly lower than that achieved by a single-layer MLP, which is consistent with the circuit complexity theory arguments put forward in Bengio et al. (2006). It also appears that convergence is not affected by the introduction of AWNN, as investigated in Figure 3 (right), which was not obvious considering the parametrization constraints encouraged by the rescaling of neurons’ activations.
@@ 5.1 Adaptation to Task Difficulty and Convergence
5.2 Training Stability Analysis
To support our argument that AWNN can reduce the time spent performing hyper-parameter selection, it is instructive to check whether AWNN learns a consistent amount of neurons across different training runs and hyper-parameter choices. Figure 4 reports the impact of the batch size and starting width averaged across the different configurations tried during model selection. Smaller batch sizes cause more instability, but in the long run we observe convergence to a similar width. Convergence with respect to different rates holds, instead, for the bounded ReLU6 activation; Appendix F shows that unbounded activations may cause the network to converge more slowly to the same width, which is in accord with the considerations about counterbalancing the rescaling effect of Section 3.1. Therefore, whenever possible, we recommend using bounded activations.
5.3 Online Network Compression via Regularization
So far, we have used an uninformative prior $p(\lambda)$ over the neural networks’ width. We demonstrate the effect of an informative prior by performing an annealing experiment on the SpiralHard dataset. We set an uninformative $p(\boldsymbol{\theta})$ and ReLU6 nonlinearity. At epoch 1000, we introduce $p(\lambda_{\ell})=\mathcal{N}(\lambda_{\ell};0.05,1)$, and gradually anneal the standard deviation up to $0.1$ at epoch $2500$. Figure 5 shows that the width of the network reduces from approximately 800 neurons to 300 without any test performance de gradation. We hypothesize that the least important neurons mostly carry negligible information, and therefore they can be safely removed without drastic changes in the output of the model. This technique might be particularly useful to compress large models with billions of parameters.
5.4 Post-hoc Trucation achieves a trade-off between performance and compute resources
To further investigate the advantages of imposing a soft ordering of importance among neurons, we show that it is possible to perform hard network’s truncation after training while still controlling the trade-off between performance and network size. Figure 6 shows an example for an MLP on the Spiral dataset, where the range of activation values (Equation 14) computed for all samples follows an exponential curve (right). Therefore, it is expected that removing the last neurons may have a negligible performance impact at the beginning and a drastic one as few neurons remain. This is what happens in the left plot, where we are able to cut an MLP with hidden width 83 by 30% without loss of accuracy, after which a smooth degradation happens. If one accepts such a trade-off, this technique may be used to distill a trained neural network at virtually zero additional cost while reducing the memory requirements.

### v3

@@ 3.1 Imposing a Soft Ordering on Neurons’ Importance
Now that the learning objective has been formalized, the missing ingredient is the definition of the neural network $p(y_{i}|\boldsymbol{\lambda}\text{=}\boldsymbol{\nu},\boldsymbol{\theta}\text{=}\boldsymbol{\rho},x_{i})$ of Equation 4 as a modified MLP. Compared to a standard MLP, we need to make use of the variational parameters $\boldsymbol{\nu}$ that affect the truncation width at each hidden layer, whereas $\boldsymbol{\rho}$ are the weights. We chose a monotonically decreasing function $f_{\ell}$, thus when a new neuron is added its relative importance is low and will not drastically impact the network output and hidden activations. In other words, we impose a soft ordering of importance among neurons.
We simply modify the classical activation $h^{\ell}_{j}$ of a hidden neuron $j$ at layer $\ell$ as
@@
where $D_{\ell-1}$ is the truncated width of the previous layer, $\sigma$ is a non-linear activation function and $w^{\ell}_{jk}\in\rho_{\ell j}$. That is, we rescale the activation of each neuron $k$ by its “importance" $f_{\ell}(j;\nu_{\ell})$. Note that the bias parameter can be part of the weight vector as usual.
It is easy to see that, in theory, the optimization algorithm could rescale the weights of the next layer by a factor $1/f_{\ell}(j;\nu_{\ell})$ to compensate for the term $f_{\ell}(j;\nu_{\ell})$. This could lead to a degenerate situation where the activations of the first neurons are small relative to the others, thus breaking the soft-ordering and potentially wasting neurons. There are two strategies to address this undesirable effect. The first is to regularize the magnitude of the weights thanks to the prior $p(\theta_{\ell+1n})$, so that it may be difficult to compensate for the least important neurons that have a high $1/f_{\ell}(j)$. The second and less obvious strategy is to prevent the units’ activations of the current layer to compensate for high values by bounding their range, e.g., using a ReLU6 activation  [51]. We apply both strategies to our experiments, although we noted that they do not seem strictly necessary in practice.
@@ 3.2 Rescaled Weight Initialization for Deep AWNN
Rescaling the activations of hidden units using Equation 10 causes activations of deeper layers to quickly decay to zero for an AWNN MLP with ReLU nonlinearity initialized using the well known Kaiming scheme  [25]. This affects convergence since gradients get close to zero and it becomes impractical to train deep AWNN MLPs. We therefore derive a rescaled Kaiming weight that guarantees that the variance of activation across layers is constant at initialization.
@@
Consider an MLP with activations as in Equation 10 and ReLU nonlinearity. At initialization, given $\alpha_{j}^{\ell}=\sigma\left(\sum_{k=1}^{D_{\ell-1}}w^{\ell}_{jk}h^{\ell-1}_{k}\right),\mathrm{Var}[w_{j*}^{\ell}]=\frac{2}{\sum_{j=1}^{D_{\ell-1}}f^{2}_{\ell}(j)}\Rightarrow\mathrm{Var}[\alpha_{j}^{\ell}]\approx\mathrm{Var}[\alpha_{j}^{\ell-1}]$.
@@
See Appendix E for an extended proof. ∎
The extended proof shows there is a connection between the variance of activations and the variance of gradients. In particular, the sufficient conditions over $\mathrm{Var}[w_{j*}^{\ell}]$ are identical if one initializes $\nu_{\ell}$ in the same way for all layers. Therefore, if we initialize weights from a Gaussian distribution with standard deviation $\frac{\sqrt{2}}{\sqrt{\sum_{j=1}^{D_{\ell-1}}f^{2}_{\ell}(j)}}$, we guarantee that at initialization the variance of the deep network’s gradients will be constant. The effect of the new initialization (dubbed “Kaiming+”) can be seen in Figure
@@
11:       $D_{\ell}\leftarrow\text{quantile function of }f_{\ell}(\cdot;\nu_{\ell})\text{\ evaluated at }k$
12:       Use $D_{\ell}$ to update $\boldsymbol{\rho}_{\ell},\boldsymbol{\rho}_{\ell+1}$ // add/remove neurons
7, where the distribution of activation values at initialization does not collapse in subsequent layers. This change drastically impacts overall convergence on the SpiralHard synthetic dataset (described in Section 4), where it appears it would be otherwise hard to converge using a standard Kaiming initialization.
3.3 Future Directions and Limitations
MLPs are ubiquitous in modern deep architectures. They are used at the very end of CNNs, in each Transformer layer, and they process messages coming from neighbors in DGNs. Our experiments focus on MLPs to showcase AWNN’s broad applicability, but there are many other scenarios where one can apply AWNN’s principles. For instance, one could impose a soft ordering of importance on CNNs’ filters at each layer, therefore learning the number of filters during training.
From a more theoretical perspective, we believe one could draw connections between our technique and the Information Bottleneck principle  [56], which seeks maximally representative (i.e., performance) and compact representations (e.g., width). Finally, the analysis of different functions $f$ to soft-order neurons, for instance a power-law distribution or a sigmoid-like function, may have a different impact on convergence and performance and could be empirically investigated.
AWNN needs to change the network at training time with an associated overhead. However, our current implementation has up to 3-4$\times$ overhead on the smallest datasets, where the inference cost is negligible compared to the network update. Although this should not be an issue with more realistic tasks, especially if one considers the time spent testing different widths during grid search, further optimizations are conceivable: i) by updating the width every $M$ mini-batches rather than at every step; ii) by allocating a larger network and dynamically grow it only if needed; iii) by creating ad-hoc support of AWNN on deep learning libraries like PyTorch.
@@ 4 Experiments and Setup
The purpose of the empirical analysis is not to claim AWNN is generally better than the fixed-width baseline. Rather, we demonstrate how AWNN overcomes the problem of fixing the number of neurons by learning it, thus reducing the amount of configurations to test. As such, due to the nature of this work and similarly to  Mitchell et al. [40], we use the remaining space to thoroughly study the behavior of AWNN, so that it becomes clear how to use it in practice. We first quantitatively verify that AWNN does not harm the performance compared to baseline models and compare the chosen width by means of grid-search model selection with the learned width of AWNN. Second, we check that AWNN chooses a larger width for harder tasks, which can be seen as increasing the hypotheses space until the neural network finds a good path to convergence. Third, we verify that convergence speed is not significantly altered by AWNN, so that the main limitation lies in the extra overhead for adapting the network at each training step. As a sanity check, we study conditions under which AWNN’s learned width does not seem to depend on starting hyper-parameters, so that their choice does not matter much. Finally, we analyze other practical advantages of training a neural network under the AWNN framework: the ability to compress information during training or post training, and the resulting trade-offs. Further analyses are in the Appendix.
@@
We compare baselines that undergo proper hyper-parameter tuning (called “Fixed”) against its AWNN version, where we replace any fixed MLP with an adaptive one. First, we train an MLP on 3 synthetic tabular tasks of increasing binary classification difficulty, namely a double moon, a spiral, and a double spiral that we call SpiralHard. A stratified hold-out split of 70% training/10% validation/20% test for risk assessment is chosen at random for these datasets. Similarly, we consider a ResNet-20  [26] trained on 3 image classification tasks, namely MNIST  [32], CIFAR10, and CIFAR100  [31], where data splits and preprocessing are taken from the original paper and AWNN is applied to the downstream classifier. In the sequential domain, we implemented a basic adaptive Recurrent Neural Network (RNN) and evaluated on the PMNIST dataset  [67] with the same data split as MNIST. In the graph domain, we train a Graph Isomorphism Network  [65] on the NCI1 and REDDIT-B classification tasks, where topological information matters, using the same split and evaluation setup of  Errica et al. [16]. Here, the first 1 hidden layer MLP as well as the one used in each graph convolutional layer are replaced by adaptive AWNN versions. On all these tasks, the metric of interest is the accuracy. Finally, for the textual domain we train a Transformer architecture  [58] on the Multi30k English-German translation task  [13], using a pretrained GPT-2 Tokenizer, and we evaluate the cross-entropy loss over the translated words. On tabular, image, and text-based tasks, an internal validation set (10%) for model selection is extracted from the union of outer training and validation sets, and the best configuration chosen according to the internal validation set is retrained 10 times on the outer train/validation/test splits, averaging test performances after an early stopping strategy on the validation set. Due to space reasons, we report datasets statistics and the hyper-parameter tried for the fixed and AWNN versions in Appendix F and G, respectively [^3] Code to reproduce results is in the supplementary material.. We ran the experiments on a server with 128 cores, 1.5TB of RAM, and 4 NVIDIA A40 with 46GB of memory.
@@ 5 Results
Table 1: Performances and total width of MLP layers for the fixed and AWNN versions of the various models used. The exact width chosen by model selection on the graph datasets is unknown since we report published results. “Linear" means the chosen downstream classifier is a linear model.
@@
We begin by discussing the quantitative results of our experiments: Table 1 reports means and standard deviations across the 10 final training runs. In terms of performance, we observe that AWNN is more stable or accurate than a fixed MLP on DoubleMoon, Spiral, and SpiralHard; all other things being equal, it seems that using more neurons and their soft ordering are the main contributing factors to these improvements. On the image datasets, performances of AWNN are comparable to those of the fixed baseline but for CIFAR100, due to an unlucky run that did not converge. There, AWNN learns a smaller total width compared to grid search.
Results on graph datasets are interesting in two respects: First, the performance on REDDIT-B is significantly improved by AWNN both in terms of average performance and stability of results; second, the total learned width is significantly higher than those tried in  Xu et al. [65], Errica et al. [16], meaning that a biased choice of possible widths had a profound influence on risk estimation for DGN models (i.e., GIN). This result makes it evident that it is important to let the network decide how many neurons are necessary to solve the task. Appendix H shows what happens when we retrain some fixed baselines using the total width as the width of each layer.
Finally, the results on the Multi30k show that the AWNN Transformer learns to use 200x parameters less than the fixed Transformer for the feed-forward networks, achieving a statistically comparable test loss. This result is appealing when read through the lenses of modern deep learning, as the power required by some neural networks such as Large Language Models  [9] is so high that cannot be afforded by most institutions, and it demands future investigations.
Adaptation to Task Difficulty and Convergence
Intuitively, one would expect that AWNN learned larger widths for more difficult tasks. This is indeed what happens on the tabular datasets (and image datasets, see Appendix I) where some tasks are clearly harder than others. Figure 2 (left) shows that, given the same starting width per layer, the learned number of neurons grows according to the task’s difficulty. It is also interesting to observe that the total width for a multi-layer MLP on SpiralHard is significantly lower than that achieved by a single-layer MLP, which is consistent with the circuit complexity theory arguments put forward in  Bengio et al. [3], Mhaskar et al. [37]. It also appears that convergence is not affected by the introduction of AWNN, as investigated in Figure 2 (right), which was not obvious considering the parametrization constraints encouraged by the rescaling of neurons’ activations.
@@
Training Stability Analysis
To support our argument that AWNN can reduce the time spent performing hyper-parameter selection, we check whether AWNN learns a consistent amount of neurons across different training runs and hyper-parameter choices. Figure 3 reports the impact of the batch size and starting width averaged across the different configurations tried during model selection. Smaller batch sizes cause more instability, but in the long run we observe convergence to a similar width. Convergence with respect to different rates holds, instead, for the bounded ReLU6 activation; Appendix J shows that unbounded activations may cause the network to converge more slowly to the same width, which is in accord with the considerations about counterbalancing the rescaling effect of Section 3.1. Therefore, whenever possible, we recommend using bounded activations.
@@
Online Network Compression via Regularization
So far, we have used an uninformative prior $p(\lambda)$ over the neural networks’ width. We demonstrate the effect of an informative prior by performing a width-annealing experiment on the SpiralHard dataset. We set an uninformative $p(\boldsymbol{\theta})$ and ReLU6 nonlinearity. At epoch 1000, we introduce $p(\lambda_{\ell})=\mathcal{N}(\lambda_{\ell};0.05,1)$, and gradually anneal the standard deviation up to $0.1$ at epoch $2500$. Figure 4 shows that the width of the network reduces from approximately 800 neurons to 300 without any test performance degradation. We hypothesize that the least important neurons carry negligible information, therefore they can be safely removed without drastic changes in the output of the model. This technique might be particularly useful to compress large models with billions of parameters.
@@
Post-hoc Truncation Achieves a Trade-off between Performance and Compute Resources
To further investigate the consequences of imposing a soft ordering among neurons, we show that it is possible to perform straightforward post-training truncation while still controlling the trade-off between performance and network size. Figure 5 shows an example for an MLP on the Spiral dataset, where the range of activation values (Equation 10) computed for all samples follows an exponential curve (right). Intuitively, removing the last neurons may have a negligible performance impact at the beginning and a drastic one as few neurons remain. This is what happens, where we are able to cut an MLP with hidden width 83 by 30% without loss of accuracy, after which a smooth degradation happens. If one accepts such a trade-off, this technique may be used to “distill” a trained neural network at virtually zero additional cost while reducing the memory requirements. Note that truncation heuristics that are either random or based on the magnitude of neurons’ activations (i.e., excluding the rescaling term) do not perform as well.